Please describe question in detail Johns Custom Computer Shop (JCCS) assembles computers for both individual and corporate customers. The company is organized
Answer #1

Part A

Personal

Business

Allocated costs

$43737

$25513

Accounts receivable cost ÷ bills prepared = 69250/950 = $72.89 per bill

Personal = 69250*600/950 = $43737

Business = 69250*350/950 = $25513

Personal

Business

Billing

21000

12250

Dispute resolution

27000

9000

Total allocated costs

$48000

$21250

Billing

Personal = (33250*600/950) = 21000

Business = (33250*350/950) = 12250

Dispute resolution

Personal = (36000*60/80) = 27000

Business = (36000*20/80) = 9000